# Planning of Off-Grid Renewable Power to Ammonia Systems with Heterogeneous Flexibility: A Multistakeholder Equilibrium Perspective

### v1

@@ IV-B Equivalent Problem for Solving the Equilibrium
[…]
@@
[…]
@@
$\partial{L}/\partial\bm{x}^{\text{rg}}=\nabla_{\bm{x}^{\text{rg}}}(\textstyle\sum_{k}C^{k})+(\bm{\lambda}^{\text{rg}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{rg}}+(\bm{\lambda}^{\text{hp}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{hp}}$
$+(\bm{\lambda}^{\text{as}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{as}}+(\rho^{\text{rg-hp,e}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
@@
[…]
where $L$ is the Lagrangian of (57); (58b) are derived similarly to (58), with the detailed expressions not given here due to space limitations; $\nabla_{\bm{x}^{\text{rg}}}(C^{\text{hp}}+C^{\text{as}})$, $\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{hp}}$, and $\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{as}}$ are all zero. As a result, (58) can be reformulated as:
$\nabla_{\bm{x}^{k}}C^{k}+(\bm{\lambda}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{g}^{k}+(\bm{\rho}^{\text{rg-hp,e}})^{\text{T}}\nabla_{\bm{x}^{k}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
@@ IV-C Temporal Decomposition-Based Solution Algorithm
$\min_{\{\bm{x}_{\text{inv}},\bm{x}_{\text{ope},l},\forall l\}}$ $\bm{a}^{\text{T}}\bm{x}_{\text{inv}}+\sum\nolimits_{l=1}^{12}\bm{b}^{\text{T}}\bm{x}_{\text{ope},l},$ (60a)
@@
$\min_{\{\bm{x}_{\text{inv}},\text{LBD}_{l},\forall l\}}~{}$ ${\bm{a}}^{\text{T}}\bm{x}_{\text{inv}}+\sum\nolimits_{l=1}^{12}\text{LBD}_{l},$
@@ V-A Case Setups
$\text{LCOA}=(\sum_{k}C^{k}-\dfrac{8760}{T}\sum_{t}M_{t}^{\text{as,sell}}\rho_{t}^{\text{a}})/(\dfrac{8760}{T}\sum_{t}M_{t}^{\text{as,pro}}).$
@@ V-B Analysis of Typical Operational and Trading Results under Equilibrium in the Base Case
Fig. 9 shows that peak, flat, and valley periods of wind and solar power correspond to low, medium, and high electricity prices under equilibrium, such as $t\in[108,117]$ h, $t\in[98,107]$ h, and $t\in[91,97]$ h, respectively. During flat periods, the prices stabilize around $0.37$ CNY/kWh. Microscopically, $\rho^{\text{rg-as,e}}$ depends more on wind and solar resources compared to $\rho^{\text{rg-hp,e}}$ due to the smaller invested capacity of the RG-side BES. During high periods, the electricity price approaches 00 CNY/kWh due to the saturated production capacities of AE and ASY, causing the renewable power supply to exceed the demand. During low periods, the prices stay around $0.6$ CNY/kWh, as it is enveloped by the backup power price $\rho^{\text{as,back}}$ ($0.6$ CNY/kWh) for the ASY.
@@
TABLE IV: Planning Results and Performances of Different Cases
@@
{$-C^{\text{rg}},-C^{\text{hp}},-C^{\text{as}}$}
@@ V-C Planning Results for the Multistakeholder Sizing Equilibrium under Different Configurations
We further analyze the MSSE under eight cases to reveal a potential dilemma in multistakeholder investment practices during implementation of ReP2A systems:
  • C1 (The base case in Section V-B): the RG and HP both invest in BES, and the HP and AS both invest in HSTs. In other words, $W^{\text{rg,bes}}$, $W^{\text{hp,bes}}$, $W^{\text{hp,hst}}$, and $W^{\text{as,hst}}$ all must be optimized, as described in Sections II and III.
@@
In case C7, where the ASY capacity is predetermined on the basis of ammonia demand, the overall technoeconomic viability of ReP2A is inferior to that in cases C1 to C6. This suggests that sizing on the basis of wind and solar resources is marginally more advantageous than sizing by ammonia demand. In C8, with all facilities’ capacities fixed, the overall technoeconomic viability of ReP2A becomes unoptimistic, and significant differences in stakeholders’ interests can be observed. Therefore, we observe that coordinated sizing for renewable, AE, and ASY facilities, although conducted by stakeholders with conflicting interests, contributes to the technoeconomic viability of ReP2A.
In summary, for cases C1 to C6, to improve the overall technoeconomic performance of the ReP2A system, the capacities of renewables, AE, ASY, and BSs must be properly aligned. However, despite efforts to optimize BES and HST configurations, stakeholders’ interests remain unbalanced due to their heterogeneous characteristics. Specifically, in no case do all three stakeholders achieve a positive profit. This phenomenon is very different from the well-studied applications in power systems, for example, games among power sources [20], microgrids [21], and prosumers [30]. Analyzing the the HP’s cost in C1 to C6, we can see that the HP plays a dominant role in the equilibrium. To make ReP2A projects invested in by multiple stakeholders implementable, benefit transfer mechanisms or pricing agreements could be established before implementing a ReP2A project to ensure the interests of all parties.
@@ V-D Recommendation for the Investment of ReP2A Projects
In cases D1 to D6, although social welfare is positive, no one achieves mutual benefits under equilibrium, which is consistent with the findings from C1 to C6. Therefore, we conclude that the multistakeholder ReP2A system can hardly achieve reasonable investment under free competition. To address this, it is recommended that regulatory authorities establish a negotiation platform for multistakeholder investment. This would allow stakeholders to form a consortium [7] and enter into benefit transfer or pricing agreements that ensure mutual benefits and facilitate successfully implementing ReP2A projects.
@@
As an example, in case D1, RG and HP are profitable, whereas AS incurs a loss. To mitigate this, we propose a benefit transfer mechanism whereby RG transfers 3% of its revenue from electricity sales to the HP and where the HP transfers 6% of its revenue from hydrogen sales to the AS. The simulation results for this scenario, denoted as E1, are recorded in Table IV, which shows that all three stakeholders achieve a positive profit. Due to the space limit, we will not discuss this further.
@@ VI Conclusions
2) To increase the attractiveness and feasibility of ReP2A projects, regulators should establish a negotiation platform where stakeholders can sign benefit transfer or pricing agreements, enabling mutual benefits and ensuring the successful implementation of ReP2A projects.
Future research will focus on the uncertainties associated with renewable resources and external ammonia prices. Additionally, the equilibrium in ammonia transactions with external chemical users within the ReP2A system will be considered.

### v2

@@ IV-B Equivalent Problem for Solving the Equilibrium
The dual variables of (17) and (34) represent the equilibrium prices for different stakeholders. When transactions are reached, trading prices w.r.t. all parties $\forall k\in\{\text{rg,hp,as}\}$ are the same, as
@@
s.t. $\bm{h}^{k}=\bm{0}{\colon\bm{\mu}^{k}},~\forall k\in\{\text{rg,hp,as}\},$
$\bm{g}^{k}\leq\bm{0}{\colon\bm{\lambda}^{k}},~\forall k\in\{\text{rg,hp,as}\},$
@@
$\partial{L}/\partial\bm{x}^{\text{rg}}=\nabla_{\bm{x}^{\text{rg}}}(\textstyle\sum_{k}C^{k})+(\bm{\lambda}^{\text{rg}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{rg}}+(\bm{\mu}^{\text{rg}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{h}^{\text{rg}}$
$+(\bm{\lambda}^{\text{hp}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{hp}}+(\bm{\mu}^{\text{hp}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{h}^{\text{hp}}$
$+(\bm{\lambda}^{\text{as}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{as}}+(\bm{\mu}^{\text{as}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}\bm{h}^{\text{as}}$
$+(\bm{\rho}^{\text{rg-hp,e}})^{\text{T}}\nabla_{\bm{x}^{\text{rg}}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
@@
$(\bm{\lambda}^{k})^{\text{T}}\bm{g}^{k}=\bm{0},~\bm{\lambda}^{k}\geq\bm{0},~\bm{g}^{k}\leq\bm{0},~\forall k\in\{\text{rg,hp,as}\},$ (59c)
$\bm{h}^{k}=\bm{0},\ \forall k$ (59d)
@@
where $L$ is the Lagrangian of (58); (59b) are derived similarly to (59), with the detailed expressions not given here due to space limitations; $\nabla_{\bm{x}^{\text{rg}}}(C^{\text{hp}}+C^{\text{as}})$, $\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{hp}}$, $\nabla_{\bm{x}^{\text{rg}}}\bm{h}^{\text{hp}}$, $\nabla_{\bm{x}^{\text{rg}}}\bm{g}^{\text{as}}$, and $\nabla_{\bm{x}^{\text{rg}}}\bm{h}^{\text{as}}$ are all zero. As a result, (59) can be reformulated as:
$\nabla_{\bm{x}^{k}}C^{k}+(\bm{\lambda}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{g}^{k}+(\bm{\mu}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{h}^{k}$
$+(\bm{\rho}^{\text{rg-hp,e}})^{\text{T}}\nabla_{\bm{x}^{k}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
@@ IV-C Temporal Decomposition-Based Solution Algorithm
$\min_{\{\bm{x}_{\text{inv}},\bm{x}_{\text{ope},l},\forall l\}}\$$\bm{a}^{\text{T}}\bm{x}_{\text{inv}}+\sum\nolimits_{l=1}^{12}\bm{b}^{\text{T}}\bm{x}_{\text{ope},l},$ (61a)
@@
$\hskip-3.0pt\min_{\{\bm{x}_{\text{inv}},\text{LBD}_{l},\forall l\}}$~ ${\bm{a}}^{\text{T}}\bm{x}_{\text{inv}}+\sum\nolimits_{l=1}^{12}\text{LBD}_{l},$
@@ V-A Case Setups
$\text{LCOA}=(\sum_{k}C^{k}-\dfrac{8760}{T}\sum_{t}M_{t}^{\text{as,sell}}\rho_{t}^{\text{a}})/(\dfrac{8760}{T}\sum_{t}M_{t}^{\text{as,pro}}).$ (70)
@@ V-B Analysis of Typical Operational and Trading Results under Equilibrium in the Base Case
Fig. 9 shows that peak, flat, and valley periods of wind and solar power correspond to low, medium, and high electricity prices under equilibrium, such as $t\in[108,117]$ h, $t\in[98,107]$ h, and $t\in[91,97]$ h, respectively. During flat periods, the prices stabilize around $0.37$ CNY/kWh. Microscopically, $\rho^{\text{rg-as,e}}$ depends more on wind and solar resources compared to $\rho^{\text{rg-hp,e}}$ due to the smaller invested capacity of the RG-side BES. During high periods, the electricity price approaches 0 CNY/kWh due to the saturated production capacities of AE and ASY, causing the renewable power supply to exceed the demand. During low periods, the prices stay around $0.6$ CNY/kWh, as it is enveloped by the backup power price $\rho^{\text{as,back}}$ ($0.6$ CNY/kWh) for the ASY.
@@
TABLE IV: Planning Results and Performance Metrics of Different Cases
@@
$C^{\text{rg}},-C^{\text{hp}},-C^{\text{as}}$}
@@ V-C Planning Results for the Multistakeholder Sizing Equilibrium under Different Configurations
We further analyze the MSSE under ten cases, examining the differences in sizing results under various configurations, to reveal a potential dilemma in multistakeholder investment practices during the implementation of ReP2A systems:
  • C1 (The base case in Section V-B): the RG and HP both invest in BESs, and the HP and AS both invest in HSTs. In other words, $W^{\text{rg,bes}}$, $W^{\text{hp,bes}}$, $W^{\text{hp,hst}}$, and $W^{\text{as,hst}}$ all must be optimized, as described in Sections II and III.
@@
  • C9: RG, HP, and AS plan separately without integration. RG generates electricity for the power grid (price = 0.18 CNY/kWh), HP buys electricity at time-of-use price [14] to produce hydrogen for the hydrogen market (selling price = 2 CNY/Nm3), and AS buys hydrogen from hydrogen market (purchase price = 2.2 CNY/Nm3) to produce ammonia.
  • C10: HP and AS are invested in by a single entity, represented as H&A, without hydrogen trading process. The electricity price between RG and H&A is defined as $\rho^{\text{rg-h\&a,e}}$. The HST is located on the ASY side and its capacity in the H&A is defined as $W^{\text{h\&a,hst}}$.
@@
In case C7, where the ASY capacity is predetermined on the basis of ammonia demand, the overall technoeconomic viability of ReP2A is inferior to that in cases C1 to C6. This suggests that sizing on the basis of wind and solar resources is marginally more advantageous than sizing by ammonia demand. It is important to note that even when the RG can optimize wind, solar, and BES capacities, its profit remains negative due to the fixed transmission line investment cost ($(1+\eta_{\text{O\&M}})\text{CRF}(r,y)C^{\text{rg,line}}$ = 15.40 million CNY). In reality, optimizing the equipment capacity generates a profit of 11.11 million CNY, resulting in a net profit of -4.29 million CNY. In C8, with all facilities’ capacities fixed, the overall technoeconomic viability of ReP2A becomes unoptimistic, and significant differences in stakeholders’ interests can be observed. Therefore, we observe that coordinated sizing for renewable, AE, and ASY facilities, although conducted by stakeholders with conflicting interests, contributes to the technoeconomic viability of ReP2A.
Under case C9, there is no electricity or hydrogen trading and delivery process among the three entities. RG does not deploy BES because the electricity price is fixed. Similarly, HP and AS do not deploy HST. HP does not deploy BES because using off-peak electricity for hydrogen production still cannot offset the BES costs. AS still configures AST to sell ammonia at higher prices in the external market. The sizing results of AE and ASY are both at their lower bounds (zero if no lower bounds exist), indicating poor technoeconomic performance and no appeal to any entity. Additionally, hydrogen production from carbon-emitting electricity for ammonia synthesis is ineligible for green subsidies [53] and lacks application potential.
A comparison of C10 with C1 and C4 shows that the overall sizing results and technoeconomic differences of the ReP2A system between the integrated H&A model and the three-entity model are minimal. However, the transactions and conflicts of interest between the entities differ significantly. The flexibility of the H&A is improved compared to the individual HP and AS, leading to a reduction in marginal production cost when participating in electricity trading. As a result, the revenue of RG is transferred to the H&A side. Therefore, encouraging companies to invest in both hydrogen and ammonia simultaneously can increase overall profits and avoid conflicts, compared to separate investments. Additionally, independent HP and AS stakeholders are incentivized to form profitable collaborations.
In summary, for cases C1 to C6, to improve the overall technoeconomic performance of the ReP2A system, the capacities of renewables, AE, ASY, and BSs must be properly aligned. However, despite efforts to optimize BES and HST configurations, stakeholders’ interests remain unbalanced due to their heterogeneous characteristics. Specifically, in no case do all three stakeholders achieve a positive profit. This phenomenon is very different from the well-studied applications in power systems, for example, games among power sources [32], microgrids [33], and prosumers [49]. Analyzing the HP’s cost in C1 to C6, we can see that the HP plays a dominant role in the equilibrium. To make ReP2A projects invested in by multiple stakeholders implementable, benefit transfer mechanisms or pricing agreements could be established before implementing a ReP2A project to ensure the interests of all parties.
@@ V-D Recommendation for the Investment of ReP2A Projects
In cases D1 to D6, although social welfare is positive, no one achieves mutual benefits under equilibrium, which is consistent with the findings from C1 to C6. Therefore, we conclude that the multistakeholder ReP2A system can hardly achieve reasonable investment under free competition. To address this, it is recommended that regulatory authorities establish a negotiation platform for multistakeholder investment. This would allow stakeholders to form a consortium [14] and enter into benefit transfer or pricing agreements that ensure mutual benefits and facilitate successfully implementing ReP2A projects. Possible methods include:
V-D1 Revenue Transfer
@@
As an example, in case D1, RG and HP are profitable, whereas AS incurs a loss. To mitigate this, we propose a benefit transfer mechanism whereby RG transfers 3% of its revenue from electricity sales (i.e., $P^{\text{rg,sell,hp}}\rho^{\text{rg-hp,e}}$) to the HP and where the HP transfers 6% of its revenue from hydrogen sales (i.e., $f^{\text{hp,sell,as}}\rho^{\text{hp-as,h}}$) to the AS. Based on the results of D1, transferring revenue from HP to AS alone cannot compensate for AS’s loss. Therefore, RG also needs to participate in the benefit transfer. The simulation results for this scenario, denoted as E1, are recorded in Table IV, which shows that all three stakeholders achieve a positive profit. Due to the space limit, we will not discuss this further.
V-D2 Profit Re-arrangement
We divide the total profit into a larger basic part and a smaller compensation part, and design a profit re-arrangement method. The basic part is arranged according to the investment cost proportions of the three parties, i.e., composition of the LCOA, ensuring that all parties receive positive base returns. The compensation part serves as an incentive for the beneficiaries (RG and HP here) under the equilibrium, and is allocated according to the proportion of their respective profits. Similarly, in Case D1, the total profit is divided according to an 8:2 ratio, with the basic part arranged approximately in a 6.0:2.7:1.3 ratio (as shown in Fig. 6) to RG/HP/AS. For example, RG’s profit is calculated as $3.40\times 0.8\times\frac{6.0}{6.0+2.7+1.3}+3.40\times 0.2\times\frac{6.34}{6.34+14.96}\approx 1.83$ million CNY. The resulting arrangement is denoted as E2, as shown in Table IV.
V-D3 Price Contacts
In addition, multiple stakeholders can sign contracts for transaction prices, making adjustments based on marginal production costs. In Case D1, RG achieves positive profits at marginal electricity prices of approximately 0.2407 CNY/kWh and 0.2635 CNY/kWh, while AS achieves positive profits at a marginal hydrogen price of around 1.810 CNY/Nm3. Based on the equilibrium prices, AS is evidently at a loss. Therefore, the hydrogen price can first be reduced based on the spot price $\rho^{\text{hp-as,h}}$ to ensure AS’s profitability. If HP’s profit becomes positive, pricing adjustments end. Otherwise, the electricity price is reduced to transfer part of RG’s operational profit to HP. Here, the three parties can sign a contract to trade at electricity and hydrogen prices $\hat{\rho}^{\text{rg-hp/as,e}}$, and $\hat{\rho}^{\text{hp-as,h}}$, which are 0.98 times and 0.944 times of the equilibrium values ($\hat{\rho}^{\text{rg-hp/as,e}}=0.98\rho^{\text{rg-hp/as,e}}$ and $\hat{\rho}^{\text{hp-as,h}}=0.944\rho^{\text{hp-as,h}}$), respectively, denoted as E3 in Table IV.
The methods above can be designed based on the application, and the parameters (such as 3%, 6%, 8:2, 0.98, and 0.944) should be tailored, aiming to create conditions for mutual benefits among all stakeholders.
V-E Computational Performance
Figure 11: Convergence and computation time of the proposed multicut and conventional BD algorithms. (a) Proposed. (b) Conventional.
[…]
This paper focuses on the planning of ReP2A but does not involve the selection of power lines, pipelines, and electrolyzers. As the number of planning objects and the dimensions of decision variables and constraints increase, long-term planning problems will become more complex, highlighting the need for more efficient decomposition algorithms.
V-F Impact of Uncertainty on Sizing Equilibrium
The capacity sizing based on typical single-year wind, solar, and price data may face issues of oversizing the system or RES curtailment due to the inter-annual variability of renewable energy [54]. Therefore, the impact of long-term uncertainties in renewable generation, and ammonia prices on the MSSE are analyzed. We divide recent years of wind and solar data into four scenarios based on renewable energy potential (i.e., FLH) [54], with the FLH shown in Fig. 12. Then, we incorporate four scenarios into the multistakeholder equilibrium framework. Since ammonia prices are negatively correlated with wind and solar resources, they are not considered separately. Furthermore, the objectives of the three stakeholders are modified as (71) and the solution approach is similarly applicable.
$C^{k}=C_{\text{inv}}^{k}+\sum\nolimits_{s=1}^{S}\pi_{s}C_{\text{ope}}^{k}(s),\ \forall k$ (71)
where $s$ and $C_{\text{ope}}^{k}(s)$ are scenario index and its operational cost, respectively; $S$ is the number of scenarios; $\pi_{s}$ represents the probability of different scenario $s$. The planning result is indicated by F1 and presented in Table IV. The algorithm converges after 106 iterations, with a computation time of 77.8 minutes. Similarly, as mentioned in Section V-E, the time becomes shorter with more abundant computational resources. Therefore, the proposed method can effectively address uncertainty through the scenario set.
Comparing F1 with C1, the capacity in F1 is slightly higher to accommodate higher renewable energy output. As renewable energy potential increases, LCOA decreases, but the drop in ammonia prices leads to a higher total cost. According to the three-stakeholder profits, the uncertainty of wind and solar power does not change the dominant role (i.e., HP) in the equilibrium, but it affects the equilibrium prices. When there is a positive forecast error in renewable energy, the electricity price decreases, and RG profit shifts to HP and AS. Hydrogen price is influenced by both electricity and ammonia prices, and the profit changes of HP and AS require further consideration.
Figure 12: FLH of renewable power under different scenarios representing variations across multiple years. (a) Wind power. (b) Solar power.
@@ VI Conclusions
2) To increase the attractiveness and feasibility of ReP2A projects, regulators should establish a negotiation platform where stakeholders can sign benefit transfer, pricing, or total profit re-arrangement agreements, enabling mutual benefits and ensuring the successful implementation of ReP2A projects.
In the future, the following areas require further work to promote the development of ReP2A, including:
1) Introducing appropriate policies to guide and improve the transaction rules for electricity, hydrogen, and ammonia.
2) Proposing long-term planning methods that consider diverse uncertainties, such as renewable generation, energy prices, and demand.
3) Developing coordinated operational mechanisms for multistakeholder ReP2A systems.
Our future research will focus on the carbon-emission market and its impact on planning. Additionally, the equilibrium in ammonia transactions with external chemical users within the ReP2A system will be considered.